OpenStudy (anonymous):

write the slope-intercept form of the equation of the line through(-4,-5), parallel to y=7x-2

OpenStudy (anonymous):

Parallel means it will have the same slope but a different y-intercept. So right now we have \[\large y=7x\] The next step is to plug in the "-4" for the x and see what you get for the y. :)

OpenStudy (anonymous):

yes so right now this is how i have it y=7x-2 -5=7(-2)-2 -5=-28-2 -5=-30 -35=b y=7x-35 I am lost from here

OpenStudy (anonymous):

Woah woah. We only have \(y=7x\) right now. We're solving from there. Can you plug -4 in for x?

OpenStudy (anonymous):

y=7(-4)-2 y=-28-2

OpenStudy (anonymous):

No -2, take that out. It would be just \(y=-28\) But it should be \(y=-5\) since the set of points is (-4,-5). So what can you add to -28 to get -5?

OpenStudy (anonymous):

23

OpenStudy (anonymous):

Right. So the equation would be \[\large y=7x+23\] and that's it, we're done.

OpenStudy (anonymous):

so can you PLEASE explain to me how i graph it?

OpenStudy (anonymous):

Manually or on a calc?

OpenStudy (anonymous):

manually on a line graph

OpenStudy (anonymous):

Just plug in an x and then mark the cooresponding y. Once you've made enough dots at points, draw a line through them. I'll give you an example for x=1 \[\large y=7(1)+23\\\large y=30\] so the point would be (1,30)